Can you please help me with these exponential/logarithmic equations?
a) solve 2^x=14
b) solve 3^x+2=81
c) solve log(x+8)= 1 + log(x-10)
Thank you so much for your help.

Answer by stanbon(75887)

a) solve 2^x=14
Take the log of both sides to get:
xlog2 = log14
x = [log14]/[log2] = 3.8074
-----------
b) solve 3^x+2=81
Rewrite as:
3^(x+2) = 3^4
x+2 = 4
x=2
-----------
c) solve log(x+8)= 1 + log(x-10)
Rewrite as:
log(x+8)-log(x-10) = 1
log[(x+8)/(x-10)] = 1
Take the anti-log to get:
(x+8)/(x-10) = 10
x-8 = 10x-100
9x= 92
x= 10.22222...
-------------------
Cheers,
Stan H.